# CasTGAN: Cascaded Generative Adversarial Network for Realistic Tabular Data Synthesis

### v1

@@ 6.4 Robustness Against Privacy Attacks
From Table 6, it can be observed that the perturbation coefficient $\epsilon$ greatly impacts the closeness of attacked samples to the training data. For unperturbed and minimally perturbed data features, it can be noticed that the attacked samples are relatively close to the training samples, indicating that the attackers might succeed in recovering training datapoints. We observe that the proximity to the training samples increases for greater $\epsilon$ values, which demonstrates the additional privacy guarantees that can be provided when altering the labels. We additionally notice how access to the data processors gives a major advantage to the attackers attempting to recover the training samples. This holds true especially for the Adult dataset, where using the trained label encodings of the auxiliary learners lead to more targeted attacks that are even closer to the training samples than attackers on unperturbed data with no access to the data preprocessors. Another interesting observation is that the attacks on the Housing dataset are greatly impacted by incremented in $\epsilon$, which is plausible, given that the dataset mainly consists of numerical features.
In addition to the proximity to the training samples, we also investigate whether perturbing the labels of the auxiliary learners can contribute to a reduction in the quality of unattacked synthetic data as demonstrated in Table 7. For the Adult, Bank and Credit datasets it can be evident that applying perturbations insignificantly impacts the the evaluation metrics of the synthetic datasets. We observe that the PR-AUC scores on the test data and KS statistic for univariate distributions are minimally influenced by the changes in $\epsilon$. In contrast, it appears that perturbing the data impacts the CODRV scores as a result of the correlation errors between the synthetic and the real datasets. Interestingly, applying perturbations on the Housing dataset appear to improve the quality of the synthetic output in addition to increasing its robustness against white-box privacy attacks.
Table 7: Impact of auxiliary learners imputation on the synthetic output.
TSTR KS CODRV
@@
7 Conclusion
@@
Given the challenging nature of generating realistic synthetic tabular data, there are several paths for future work. This work can be extended by incorporating additional data types within the tabular data generation such as free text and timestamps. Moreover, we intend to explore how our approach can generate more diversified combination of categories, while maintaining its ability in minimising the number of invalid data records.
Acknowledgments
This project was supported by DNB ASA through the funding of this research.
Appendix A Training Details
A.1 Hyperparameters
Table 8: Hyperparameters used for all datasets
@@ A.1 Hyperparameters
A.2 Losses
@@
Figure 3: Training losses
A.3 Training Elapsed Duration
Figure 4: Comparison of run times between the different baselines
Appendix B Further Results
B.1 TSTR
Table 9: Adult Dataset Classification
Table 10: Bank Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
train 0.8550 0.9075 0.6663 0.7744
tablegan 0.5617 0.4102 0.1376 0.2225
medgan 0.7920 0.8038 0.4888 0.5777
ctgan 0.8304 0.8751 0.6235 0.6932
cWGAN 0.7586 0.5887 0.0002 0.3030
ctab-gan 0.8400 0.8849 0.6542 0.7192
CasTGAN 0.8272 0.8779 0.6380 0.6718
Accuracy ROCAUC F1-score PR-AUC
train 0.9084 0.9342 0.4978 0.6085
tablegan 0.2149 0.3910 0.1670 0.0954
medgan 0.8803 0.6492 0.1547 0.2729
ctgan 0.8979 0.8474 0.3492 0.4805
cWGAN 0.8337 0.6864 0.1297 0.2692
ctab-gan 0.8970 0.8768 0.3550 0.4920
CasTGAN 0.8991 0.9187 0.5823 0.5657
Table 10: Bank Dataset Classification
Table 11: Credit Dataset Classification
Table 12: Diabetes Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
train 0.8201 0.7685 0.4663 0.5374
tablegan 0.2185 0.4605 0.3482 0.2084
medgan 0.6807 0.6146 0.3448 0.3050
ctgan 0.7826 0.6845 0.4561 0.4558
cWGAN 0.7760 0.5253 0.0053 0.2166
ctab-gan 0.8138 0.7136 0.4708 0.4911
CasTGAN 0.8048 0.7315 0.2845 0.4995
Accuracy ROCAUC F1-score PR-AUC
train 0.8623 0.8142 0.2545 0.3980
tablegan 0.8604 0.5222 0.0000 0.1585
medgan 0.8519 0.7544 0.2473 0.3241
ctgan 0.8333 0.7990 0.3924 0.3711
cWGAN 0.8443 0.7636 0.2749 0.3325
ctab-gan 0.8432 0.7857 0.3519 0.3589
CasTGAN 0.8557 0.8077 0.3336 0.3866
Table 12: Diabetes Dataset Classification
Table 13: Cars Dataset Regression
Table 14: Housing Dataset Regression
RMSE R2 Score
train 4330.3428 0.7604
tablegan 16737.5294 -2.0148
medgan 4822.8477 0.7491
ctgan 6558.5413 0.5327
cWGAN 166818.8351 -319.0963
CasTGAN 6396.3703 0.5566
RMSE R2 Score
train 393507.3597 -0.3696
tablegan 2943381.0066 -83.1326
medgan 405452.4482 -0.3845
ctgan 458694.4080 -0.6533
cWGAN 518963.3807 -1.1054
ctab-gan 472485.7269 -0.7284
CasTGAN 419848.1947 -0.4433
Table 14: Housing Dataset Regression
B.2 Statistical Similarity
(a) Categorical features
(b) Numerical features
Figure 5: Adult dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 6: Bank dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 7: Cars dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 8: Diabetes dataset statistical similarity

### v2

@@ VI-D Robustness Against Privacy Attacks
From Table VII, it can be observed that the perturbation coefficient $\epsilon$ greatly impacts the closeness of attacked samples to the training data. For unperturbed and minimally perturbed data features, it can be noticed that the attacked samples are relatively close to the training samples, indicating that the attackers might succeed in recovering training datapoints. We observe that the proximity to the training samples increases for greater $\epsilon$ values, which demonstrates the additional privacy guarantees that can be provided when altering the labels. We additionally notice how access to the data processors gives a major advantage to the attackers attempting to recover the training samples. This holds true especially for the Adult dataset, where using the trained label encodings of the auxiliary learners lead to more targeted attacks that are even closer to the training samples than attackers on unperturbed data with no access to the data preprocessors. Another interesting observation is that the attacks on the Housing dataset are greatly impacted by incremented in $\epsilon$, which is plausible, given that the dataset mainly consists of numerical features.
In addition to the proximity to the training samples, we also investigate whether perturbing the labels of the auxiliary learners can contribute to a reduction in the quality of unattacked synthetic data as demonstrated in Table VIII. For the Adult, Bank and Credit datasets it can be evident that applying perturbations insignificantly impacts the the evaluation metrics of the synthetic datasets. We observe that the PR-AUC scores on the test data and KS statistic for univariate distributions are minimally influenced by the changes in $\epsilon$. In contrast, it appears that perturbing the data impacts the CORDV scores as a result of the correlation errors between the synthetic and the real datasets. Interestingly, applying perturbations on the Housing dataset appear to improve the quality of the synthetic output in addition to increasing its robustness against white-box privacy attacks.
TABLE VIII: Impact of auxiliary learners imputation on the synthetic output.
TSTR KS CORDV
@@
VI-E Impact of Auxiliary Learners Loss
We further analyse the impact of the auxiliary learners on the quality of the synthetic data samples produced by CasTGAN. To this end, we conduct experiments on our model by tuning the auxiliary loss coefficient parameters $\lambda_{AL_{1}},\lambda_{AL_{2}},\ldots,\lambda_{AL_{M}}$. This is implemented by adjusting $\lambda_{AL_{1}}$ and $\lambda_{AL_{M}}$, as the auxiliary loss coefficients in between are linearly and equidistantly scaled between the loss coefficients of the first and final auxiliary learners. Given that there is an indefinite number of auxiliary loss coefficients combinations and a diverse set of evaluation metrics and datasets, we conduct our the analysis on the Adult dataset by experimenting with a set of auxiliary loss coefficient values: $\{0,0.25,0.5,0.75,1\}$. As the focus of our work is improving the realistic-ness and maximising the number of semantically valid synthetic records, we quantify the number of invalid records of the synthetic samples as per the description of invalid records of the Adult dataset defined in Section VI-C. Figure 6 demonstrates the fraction of invalid records generated by CasTGAN for various auxiliary loss coefficients.
Figure 6: Proportion of invalid synthetic records of the Adult dataset generated by CasTGAN using various auxiliary loss coefficient settings.
As shown in Figure 6, adjusting the auxiliary loss coefficients is reflected by a modest change in the ratio of invalid synthetic records of the Adult dataset. As demonstrated, it appears that using a fixed auxiliary loss coefficient across the auxiliary learners by setting the value $\lambda_{AL_{1}}$ and $\lambda_{AL_{M}}$ generally increases the number of invalid records. Meanwhile, it can be observed that having a smaller difference between $\lambda_{AL_{1}}$ and $\lambda_{AL_{M}}$ leads to an improved performance in contrast to when the difference is $0.5$ or greater in general. We can notice that the lowest fraction of invalid synthetic observations is obtained by setting $\lambda_{AL_{1}}=0.25$ and $\lambda_{AL_{M}}=0$. By comparing the ratio of invalid records in Table VI and Figure 6, it is clear that the default auxiliary loss coefficient parameters that we used in our experimental setup do not yield the most optimal results, whereas further hyperparameter tuning can in practice help reduce the number of invalid records. Meanwhile, it is evident that not relying on auxiliary learners by setting the loss coefficients to zero contributes to the highest fraction of invalid synthetic records. However, we can observe that the ratio of invalid records generated when using loss coefficients of 00, thereby nullifying the effect of auxiliary learners, is nevertheless lower than the proportion of the invalid synthetic records generated by the baseline models, as displayed in Table VI. This suggests that while the improved quality of synthetic records can be to a great extent be attributed to the design of the cascaded generator architecture and employment of the WGAN-GP, the use of auxiliary learners and fine-tuning the auxiliary loss coefficients can further contribute to the generation more semantically valid records.
VII Discussion
Data in tabular form is widely used across organisations in various domains for decision support systems. In such systems, data mining techniques and statistical models are employed to enhance the predictability of significant events and to gain a deeper insight into user behaviour. Moreover, mixed-type tabular data is commonly used for facilitating knowledge exchange between domain experts and stakeholders. Given the privacy concerns associated with sharing confidential information, in addition to the scarcity of available data to share across organisations, there has been a growing interest in tabular data synthesis using generative adversarial networks, in which realistic synthetic data can be safely shared across operational units. In literature, the focus of generative models has been primarily directed towards generating synthetic data that yields acceptable machine learning utility and exhibits statistical properties which are similar to the real data. While these objectives are relatively significant, an underlying limitation is that the semantic integrity of the synthetic output is not thoroughly examined. Given that GANs attempt to approximate the data distributions loss minimising functions for the generator and the discriminator, it is imperative that GAN-based synthesis models are unable explicitly distinguish between semantically valid and invalid data records.
To this end, the cascaded tabular GAN architecture we propose in this study contributes to the reduction of the number of invalid generated observations by dedicating a generator for every feature of the dataset. The sequential cascading of generator passes incomplete synthesisation of the data to subsequent generators, in which each generator attempts to predict its target feature from the incomplete feature space it receives from the preceding generator. As the generator fills the remaining feature space with noise, the output is queried against the discriminator and the discriminator loss for the queried sample is backpropagated to the generator for improving its overall generation. For obtaining meaningful feedback to the generator’s designated feature, an auxiliary learner is coupled to the generator that calculates the prediction loss for the specific feature and propagates the loss to the generator, such that it can learn to reduce the auxiliary learner’s loss during GAN training. In such an architecture, the GAN attempts to learn the dependencies between the dataset’s features, and thereby gradually reduces the losses for the components of the GAN during the training process. Hence, the aim of CasTGAN is to increase the semantic integrity of the synthetic and reduce the invalid records that are generated. This is highly significant as the generation of invalid records may have adverse effects on knowledge sharing, potentially fostering an inaccurate comprehension of the data among stakeholders and other external entities granted access to the synthesized data.
Notably, our results indicate where our proposed generative model has remarkably performed. First, the machine learning utility results demonstrate that the synthetic output from CasTGAN is particularly useful for constructing predictive models, as reflected by the classification and regression performance metrics. The results show that CasTGAN is competitive with the state-of-the-art tabular GANs, outperforming them on some datasets, and barely falls short to the predictive performance obtained by training on the real data. Regarding the univariate variable properties, we visually demonstrate the synthetic output has a striking similarity to the the real data numerical and categorical variables. Furthermore, it can be observed that quantitative similarity analysis indicates a very close similarity between the data distributions of the real and the CasTGAN synthetic output. In accordance to addressing the gap in literature for better exploring the semantic integrity of the synthetic data, we evaluate the output validity of our synthetic data on several fronts. The correlation mapping shows that the synthetic data from our model exhibits a strong resemblance to the real data used for training. We also notice that CORDV metric which we design for measuring the correlation error as a fraction of the unique pairwise combination successfully shows that our approach is well capable of capturing the dependencies between the data features. In general, the performance of CasTGAN for machine learning usability, statistical similarity and correlation and diversity was mostly comparable with CTGAN and CTAB-GAN. This can be attributed to the use of WGAN-GP in the case of CTGAN, or classification loss and information loss in the case of CTAB-GAN, both of which contribute to the training stability and the mitigation against mode collapse. Meanwhile, calculating the ratio of invalid records on a number of datasets demonstrates that CasTGAN excels in reducing the number of invalid data observations in comparison to the existing tabular GAN approaches, thereby improving the potential of using the synthetic data for knowledge exchange. Further analysis unveils that applying perturbations on the auxiliary learners can increase the robustness of our model against privacy attacks without notably sacrificing our model’s synthetic output quality.
We also note that using the auxiliary learners leads to a more conservative approach for the GAN training process. The reduction of the invalid records comes at the expense of reducing the number of unique pairwise combinations of the data categories that CasTGAN can synthesise, and additionally not fully simulating numerical values around the boundaries for numerical features. Nevertheless, we highlight that the reduction of invalid records and improving the model’s representation of feature dependency was the focus of this work, and that CasTGAN demonstrates success in these two aspects.
VIII Conclusion
@@
Given the challenging nature of generating realistic synthetic tabular data, there are several paths for future work. This work can be extended by incorporating additional data types within the tabular data generation such as free text and timestamps. Due to cascaded architecture and the presence of multiple auxiliary learners, we point out that our framework does not offer improvements to the training speed over the existing tabular GAN models, which presents a potential opportunity for future optimisation efforts. Moreover, we intend to explore how our approach can generate more diversified combination of categories, while maintaining its ability in minimising the number of invalid data records.
Appendix A Implementation Details
A-A Hyperparameters
TABLE IX: CasTGAN hyperparameters used for all datasets.
@@ A-A Hyperparameters
A-B Strategies for Hyperparameter Configuration
There are various hyperparameters in CasTGAN, and therefore selecting the optimal hyperparameters is a non-trivial task. Defining what is considered as optimal also poses a challenge, as there may be a trade-off between enhancing the suitability of synthetic output for machine learning model fitting and increasing the semantic validity of the synthetic data. In light of this, we outline the parameters that have some significant impact on training our GAN model, and we describe how they can be tuned for optimising the performance. The parameters with less significance in correspondence to the quality of the synthetic data can utilise the default values from Appendix A-A.
Epochs: Tuning the number of training iterations can offer improvements to the quality of the synthetic output. For smaller datasets, it can be sufficient to train for a smaller number of iterations, in which the training completes faster. Meanwhile, larger datasets may require a greater number of epochs to demonstrate an improved synthetic output. It is recommended to train for a minimum of 100 epochs, though training the GAN for too long can lead to the memorisation of records by the GAN, inducing potential data records leakage and susceptibility to privacy attacks.
Batch size: Batch size refers to the number of samples stored in memory for updating the internal model parameters at every training iteration. A larger batch size is more suitable for datasets with a greater number of features and categories, whereas datasets with fewer features can be sufficiently trained with a smaller batch size. It is worth noting that a larger batch size contributes to a relatively slower training of the GAN and can contribute to a poor model convergence if not trained for a sufficient number of epochs.
Generators noise input dimension: The optimal size of the noise vector used as the input for the generators is best determined through experimentation. Wasserstein GANs commonly exhibit a higher resolution of the synthetic output through the utilisation of larger noise vector [54], though the significance of input noise size is largely dependent on the data and the other model parameters.
Generators primary networks hidden sizes: The number and the sizes of the hidden layers can significantly impact the quality of the synthetic output from the generators. Despite designating a generator for each data feature, the generators employ a common hidden layer configuration rather than a specific configuration for each generator. As each generator is responsible for generating one primary feature, the linear layers do not need to be overly sophisticated.
Discriminator primary networks hidden sizes: The architecture of the discriminator should be able to foster the competitiveness between the discriminator and the generator, hence maintain the adversary throughout the GAN training. If the neural network architecture is relatively simpler than that of the generators, the discriminator will poorly distinguish between the synthetic and the real samples. Meanwhile, a complex discriminator architecture might suppress the learning of the generators, thus, degrade the synthetic output quality of the generators. Therefore, the number and the size of the discriminator’s hidden layers should be comparable to those employed by the generators.
Discriminator updates per generators update: It might be worthwhile to adjust the number of discriminator iterations for each generator iteration, if it helps in converging the losses during training. This might be particularly the case if the discriminator is visibly weaker than the generator. If there is no major disparity between the discriminator’s architecture and the generators’ hidden layers, then there is no proven benefit in employing more than 1 discriminator gradients update for each step of the generators training [55].
Wasserstein loss gradient penalty ($\lambda_{GP}$): We find that using a Wasserstein loss gradient penalty of 10 works well, based on the training stability during our preliminary model conceptualisation. This is supported by the authors of the original WGAN-GP paper [13], who found the gradient penalty of 10 works well on various datasets and GAN architectures. This is also in accordance to the baseline tabular GANs with WGAN-GP we compare our approach against [35] [28], whom also employ $\lambda_{GP}=10$.
Auxiliary learner 1 loss coefficient ($\lambda_{AL_{1}}$): The first auxiliary learner in the cascaded generator is assigned with the largest loss coefficient among to ensure that the univariate statistical representation of the first features of the data conforms to the real data distributions. The particular value for the first auxiliary learner’s coefficient is chosen such that large prediction errors attributed to the deviation in distributions are sufficiently penalised, without making it too large such that training instability is introduced to the generators.
Auxiliary learner M loss coefficient ($\lambda_{AL_{M}}$): The final auxiliary learner can benefit from low loss coefficient values to minimise the errors propagated to the preceding generators. The purpose of the auxiliary loss coefficient of auxiliary learner M is to reduce the correlation error in the synthetic data towards the end of cascaded generator, as opposed to the first generators that prioritise the univariate similarity. In a similar manner to the first auxiliary learner, this is a hyperparameter that is largely dependent on the dataset used for training and sampling of CasTGAN. It is however advisable to set $\lambda_{AL_{M}}$ with a very small value if a small value is selected for $\lambda_{AL_{1}}$.
A-C Train on Synthetic, Test on Real
TABLE X: Adult Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
Identity 0.8550 0.9075 0.6663 0.7744
table-GAN 0.5617 0.4102 0.1376 0.2225
medGAN 0.7920 0.8038 0.4888 0.5777
CTGAN 0.8304 0.8751 0.6235 0.6932
cWGAN 0.7586 0.5887 0.0002 0.3030
CTAB-GAN 0.8400 0.8849 0.6542 0.7192
CasTGAN 0.8272 0.8779 0.6380 0.6718
TABLE XI: Bank Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
Identity 0.9084 0.9342 0.4978 0.6085
table-GAN 0.2149 0.3910 0.1670 0.0954
medGAN 0.8803 0.6492 0.1547 0.2729
CTGAN 0.8979 0.8474 0.3492 0.4805
cWGAN 0.8337 0.6864 0.1297 0.2692
CTAB-GAN 0.8970 0.8768 0.3550 0.4920
CasTGAN 0.8991 0.9187 0.5823 0.5657
TABLE XII: Credit Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
Identity 0.8201 0.7685 0.4663 0.5374
table-GAN 0.2185 0.4605 0.3482 0.2084
medGAN 0.6807 0.6146 0.3448 0.3050
CTGAN 0.7826 0.6845 0.4561 0.4558
cWGAN 0.7760 0.5253 0.0053 0.2166
CTAB-GAN 0.8138 0.7136 0.4708 0.4911
CasTGAN 0.8048 0.7315 0.2845 0.4995
TABLE XIII: Diabetes Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
Identity 0.8623 0.8142 0.2545 0.3980
table-GAN 0.8604 0.5222 0.0000 0.1585
medGAN 0.8519 0.7544 0.2473 0.3241
CTGAN 0.8333 0.7990 0.3924 0.3711
cWGAN 0.8443 0.7636 0.2749 0.3325
CTAB-GAN 0.8432 0.7857 0.3519 0.3589
CasTGAN 0.8557 0.8077 0.3336 0.3866
TABLE XIV: Cars Dataset Regression
RMSE $R^{2}$ Score
Identity 4330.3428 0.7604
table-GAN 16737.5294 -2.0148
medGAN 4822.8477 0.7491
CTGAN 6558.5413 0.5327
cWGAN 166818.8351 -319.0963
CasTGAN 6396.3703 0.5566
TABLE XV: Housing Dataset Regression
RMSE $R^{2}$ Score
Identity 393507.3597 -0.3696
table-GAN 2943381.0066 -83.1326
medGAN 405452.4482 -0.3845
CTGAN 458694.4080 -0.6533
cWGAN 518963.3807 -1.1054
CTAB-GAN 472485.7269 -0.7284
CasTGAN 419848.1947 -0.4433
A-D CasTGAN Training Stability
@@
Figure 7: CasTGAN training loss plots.